DRESS Inc. is one of the newest clothing manufacturers located in Laguna. Their product lines include sandos, round neck shirts, fitted pants, v-neck shirts, wrap around skirt, and straight cur denims. Lately, the company has been experiencing a large number of rejects during the quality inspection at the finishing line. As the head of QC Department of the company, you were asked by the President to solve this alarming QC problem. The following are the data collected in the finishing section of the company. After inspection, a number of nonconformities were found in the production of 200 sandos. SUBGROUP # of loose threads a. Determine the [1] centerline, [2] upper control limit and [3] lower control limit. 1 2 3 45 6 7 8 9 10 11 12 13 14 5 5 6 7 10 6 5 12 13 7 6 4 2 10 b. does this production of sandos appear to exhibit statistical control with respect to the quality characteristics examined in this inspection?[4] Why?15]
Answer #1

Subgroup | Sample (n) | # of loose threads | Fraction detectives 200 200 200 200 200 200 200 200 200 200 200 200 200 200 2800 0.025 0.025 0.030 0.035 0.050 0.030 0.025 0.060 0.065 0.035 0.030 0.020 0.010 0.050 2 4 10 12 13 10 12 13 14 Total 4 2 10 98 2nt 280o 14 98 Central line-C1-p--=-= 0.035 mi 2800Upper Control Limit (UCL) =p+3E1-p = 0.035 +3 10.035 (1-0.035) -0.035 +3(0.012995) 0.035 + 0.038986 UCL-0.073986 Lower Control Limit (LCL)- p-3, -035 (1-0.035) 200 = 0.035-3 =0.035-3(0012995) =0.035-0.038986 LCL=-0.00399 → 0 When the lower control limit id negative then it is rounded up to zero.0.080 0.070 0.060 0.050 defectives 0.040 CL 0.030 0.020 LCL 0.010 0.000 1 23 4 5 67 8 9 10 11 12 13 14 (b) This production pr